# Kernel Ridge Regression with Predicted Feature Inputs and Applications to Factor-Based Nonparametric Regression

### v1

@@ B.3.4 Relating $\widehat{R}(\delta)$ to $R(\delta)$
The operator $T_{K}$ is defined in the same way as the integral operator $L_{K}$ in (20) except that their domains and ranges are distinct. It is worth mentioning that $L_{K}$ and $T_{K}$ share the same eigenvalues $\{\mu_{j}\}_{j\geq 1}$, as discussed in Caponnetto and De Vito (2007). Recall the definition of $\widehat{D}(\delta)$ from (B.3.3). Since $\widehat{\mu}_{j}=0$ for all $j>n$, for any $\delta>0$, we have
@@
$\widehat{D}(\delta)\leq C\left(D(\delta)+\frac{\kappa^{2}}{n\delta}\log\frac{2}{\eta}\right)$
@@
To verify the statements, by using (13) and (14), we first note that
@@
with $[\mathbf{K}_{x}]_{ij}=n^{-1}K(\widehat{g}(X_{i}),\widehat{g}(X_{j}))$ for $i,j\in[n]$. For $K$ that is invariant to orthogonal transformations, we have
@@
for all fixed $Q\in\mathbb{O}_{r\times r}$. Therefore, repeating the arguments in (17) and (3.1) gives
@@ B.7 Proof of Corollary 3
[…]
@@ B.8 Proof of Corollary 4
[…]
@@
B.9 Proof of Theorem 3: the minimax lower bound
@@
[…]
@@
Since $\vartheta$ can be arbitrarily small and the leftmost side of (B.9) is independent of $\vartheta$, we obtain
@@
B.10 Proofs of Section 5: general convex loss functions
@@
$\mathcal{E}(f\circ\widehat{g})=$ $\mathbb{E}\left[L(Y,(f\circ\widehat{g})(X))-L(Y,f^{*}(Z))\right]$
@@
[…]
@@ B.10.1 Proof of Theorem 4
This result replaces (B.2) in the proof of Theorem 2. For Case (1) considered in the proof of Theorem 2, the $\widehat{\alpha}$ defined therein exists due to the convexity of $L(y,\cdot)$. So the proof follows analogously. For Case (2), the chain of inequalities in (B.2) can be replaced by
@@
$\leq 2C_{U}\|f_{\mathcal{H}}\|_{K}^{2}~{}\mathbb{E}\Delta_{\widehat{g}}+{2C_{U}\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}}$ by (3.1).
@@
B.10.2 Uniform upper bounds of $\mathbb{E}[\ell_{f}(Y,X)]$ over $f\in\mathcal{F}_{b}$
@@
so that the function class $\mathcal{F}_{\ell}$ is uniformly bounded within $[-3C_{\ell}\kappa\|f_{\mathcal{H}}\|_{K},3C_{\ell}\kappa\|f_{\mathcal{H}}\|_{K}]$.
@@
[…]
which implies that $\ell_{f}(Y_{i},X_{i})$ is $(C_{\ell})$-Lipschitz in $\xi_{f}(X_{i})$ for all $i\in[n]$. Then, by repeating the arguments of proving (B.2.1), and together with (179), we conclude
@@
Finally, the proof is completed by invoking Lemma 22 with $-a=b=3C_{\ell}\kappa\|f_{\mathcal{H}}\|_{K},B=V$, $L=4C_{U}\left(\|f_{\mathcal{H}}\|_{K}^{2}\mathbb{E}\Delta_{\widehat{g}}+\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}\right)$ and $\theta=2$. ∎
@@
Appendix C Auxilliary Lemmas
@@
C.1 Definition and properties of the sub-root functions
@@
If $\psi:[0,\infty)\to[0,\infty)$ is a nontrivial [^3] Functions other than the constant function $\psi\equiv 0$. sub-root function, then it is continuous on $[0,\infty)$ and the equation $\psi(\delta)=\delta$ has a unique positive solution $\delta_{\star}$. Moreover, for all $\delta>0,\delta\geq\psi(\delta)$ if and only if $\delta_{\star}\leq\delta$. Here, $\delta_{\star}$ is referred to as the fixed point.
@@
C.2 Local Rademacher complexity
@@
C.3 Auxiliary concentration inequalities

### v2

@@ A.3.4 Relating $\widehat{R}(\delta)$ to $R(\delta)$
The operator $T_{K}$ is defined in the same way as the integral operator $L_{K}$ in (19) except that their domains and ranges are distinct. It is worth mentioning that $L_{K}$ and $T_{K}$ share the same eigenvalues $\{\mu_{j}\}_{j\geq 1}$, as discussed in Caponnetto and De Vito (2007). Recall the definition of $\widehat{D}(\delta)$ from (A.3.3). Since $\widehat{\mu}_{j}=0$ for all $j>n$, for any $\delta>0$, we have
@@
$\widehat{D}(\delta)\leq C\left(D(\delta)+\frac{\kappa^{2}}{n\delta}\log\frac{2}{\eta}\right).$
@@
To verify the statements, by using the closed-form solution of KRR in (13) and (14), we first note that
@@
with $\mathbf{K}_{x}]_{ij}=n^{-1}K(\widehat{g}(X_{i}),\widehat{g}(X_{j}))$ for $i,j\in[n]$. For $K$ that is invariant to orthogonal transformations, we have
@@
for all fixed $Q\in\mathbb{O}_{r\times r}$. Therefore, repeating the arguments in (16) and (17) gives
@@ A.7 Proof of Corollary 3
$\leq\frac{C^{\prime}}{n}\delta^{\frac{2\alpha-1}{2\alpha}}+\frac{C}{n}\int_{d(\delta)+1}^{\infty}t^{-2\alpha}dt$
@@ A.8 Proof of Corollary 4
$R^{2}(\delta)$ $~\leq~\frac{\delta}{n}(d(\delta)+1)+\frac{1}{n}\int_{d(\delta)+1}^{\infty}\exp(-\gamma t)dt$
@@
A.9 Proof of Theorem 2: the minimax lower bound
@@
Let $\text{KL}(\cdot~\|~\cdot)$ denote the KL divergence between two distributions. For any $(i,j)\in[m]\times[m]$, observe that
@@
Since $\vartheta$ can be arbitrarily small and the leftmost side of (A.9) is independent of $\vartheta$, we obtain
@@
A.10 Proofs of Section 5: general convex loss functions
@@
$=$~ $\mathbb{E}\left[\ell_{f}(Y,X)\right]+\mathbb{E}\left[L(Y,(f_{\mathcal{H}}\circ\widehat{g})(X))-L(Y,f^{*}(Z))\right]$
$\leq$~ $\mathbb{E}\left[\ell_{f}(Y,X)\right]+C_{U}\mathbb{E}\left[(f_{\mathcal{H}}\circ\widehat{g})(X)-f^{*}(Z)\right]^{2}$ by 9
@@ A.10.1 Proof of Theorem 3
This result replaces (A.2) in the proof of Theorem 1. For Case (1) considered in the proof of Theorem 1, the $\widehat{\alpha}$ defined therein exists due to the convexity of $L(y,\cdot)$. So the proof follows analogously. For Case (2), the chain of inequalities in (A.2) can be replaced by
@@
$\leq 2C_{U}\|f_{\mathcal{H}}\|_{K}^{2}~\mathbb{E}\Delta_{\widehat{g}}+{2C_{U}\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}}$ by (17).
@@
A.10.2 Uniform upper bounds of $\mathbb{E}[\ell_{f}(Y,X)]$ over $f\in\mathcal{F}_{b}$
@@
so that the function class $\mathcal{F}_{\ell}$ is uniformly bounded within $3C_{\ell}\kappa\|f_{\mathcal{H}}\|_{K},3C_{\ell}\kappa\|f_{\mathcal{H}}\|_{K}]$.
@@
$\leq$~ $C_{\ell}~\bigl|\xi_{f_{1}}(x)-\xi_{f_{2}}(x)\bigr|$
which implies that $\ell_{f}(Y_{i},X_{i})$ is $(C_{\ell})$-Lipschitz in $\xi_{f}(X_{i})$ for all $i\in[n]$. Then, by repeating the arguments of proving (A.2.1), and together with (172), we conclude
@@
Finally, the proof is completed by invoking Lemma 22 with $a=b=3C_{\ell}\kappa\|f_{\mathcal{H}}\|_{K},B=V$, $L=4C_{U}\left(\|f_{\mathcal{H}}\|_{K}^{2}\mathbb{E}\Delta_{\widehat{g}}+\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}\right)$ and $\theta=2$. ∎
@@
Appendix B Auxilliary Lemmas
@@
B.1 Definition and properties of the sub-root functions
@@
If $\psi:[0,\infty)\to[0,\infty)$ is a nontrivial [^3] Functions other than the constant function $\psi\equiv 0$. sub-root function, then it is continuous on $0,\infty)$ and the equation $\psi(\delta)=\delta$ has a unique positive solution $\delta_{\star}$. Moreover, for all $\delta>0,\delta\geq\psi(\delta)$ if and only if $\delta_{\star}\leq\delta$. Here, $\delta_{\star}$ is referred to as the fixed point.
@@
B.2 Local Rademacher complexity
@@
B.3 Auxiliary concentration inequalities